# On the Utility of Accounting for Human Beliefs about AI Intention in Human-AI Collaboration

### v2

@@ 3.3. Designing Explicable AI Policy
To train an explicable AI policy, we incorporate the models of human beliefs into the AI training process. Instead of solely aiming to maximize rewards, we also focus on aligning the AI policy to increase the probability that humans can infer AI intentions based on AI actions, leveraging the human belief model to account for the human inference process. More specifically, we utilize self-play to train models. In addition to the original MDP rewards, our approach includes additional bonus rewards when the goal inference based on the human belief model is consistent with the specified goal. Specifically, we set the bonus reward to be proportional to the log likelihood of the belief model inference, $\log(\pi(a|s,g))$. With this setting, the trained AI policy would aim to reach the goal while also maximizing the likelihood of revealing its goals, where the likelihood of the goal inference follows the belief model.
3.4. Designing Collaborative AI Agents
We now describe how we design collaborative AI agents that account for different models of human behavior and beliefs. Later, in the experiments, we illustrate the effectiveness of incorporating models of human beliefs by examining the performance of AI agents paired with humans in human-AI collaboration through simulations and real-world human-subject experiments.
@@ 3.4. Designing Collaborative AI Agents
Collaborative AI agents. Using the methodology above, we have designed several AI agents based on different assumptions of the human counterpart.
@@
    Assuming humans are optimal. We first train an AI agent that learns to collaborate with itself through self-play. This is equivalent to assuming the human counterpart is acting optimally.
@@
    Assuming humans are behavioral level-00 agents. We next train an AI agent that assumes the human is a behavioral level-00 agent, using behavioral cloning to train the behavioral model.
@@
In addition to the AI agents trained using our methodologies, we also include two state-of-the-art baselines, Fictitious Co-Play (Strouse et al., 2021) and Maximum Entropy Population-based Training (Zhao et al., 2023), for comparison in our experiments. Both algorithms develop a single best response model in response to a population agents with diverse behaviors, thereby enhancing the robustness and generality of the learned best response model. The details of the setup and implementations are provided in the appendix.
4. Experiments
@@
To evaluate our approaches, we have conducted multiple sets of human-subject experiments through recruiting close to 1,000 participants from Amazon Mechanical Turk (MTurk). The experiments are approved by the IRB of our institution. Workers were paid \$1 base payments with the potential for bonus payments in some experiments. The average hourly rate is approximately \$$\$14$ across all our experiments.
4.1. Experiments Overview
@@
The main purpose of our experiments is to demonstrate that incorporating human beliefs about AI intentions into the design of AI agents can significantly improve the performance of human-AI collaboration. Specifically, we conducted three sets of experiments, beginning with examining the efficacy of our models of human behavior and beliefs, and progressing to evaluate the performance improvements.
In the first experiment, we examine the effectiveness of using behavioral cloning as a model to mimic human behavior. We recruit real-world participants to play against a predefined AI agent to collect their behavioral traces and then evaluate the prediction accuracy of using behavioral cloning to model human behavior. In the second set of experiments, we examine whether we can leverage the model of human beliefs to design an explicable AI policy (as described in Section 3.3). Specifically, we recruit real-world participants, provide them with behavioral traces generated by different AI policies, and evaluate whether the AI policies that account for human belief models indeed produce behavioral traces that make it easier for humans to infer AI intentions. Finally, in the third experiment, we examine whether the collaborative AI that accounts for human beliefs lead to higher performance in human-AI collaboration. Again we recruit real-world human participants, pair them with different AI agents (as described in Section 3.4), and measure the overall team performance.
4.2. Experiment Environment
@@
Our experiments are conducted in a grid world environment with two players and multiple goals. [^3] We conducted experiments in two variations of the environment. In the first variation, we designed the environments such that the two players are not in the same grid world. However, each player has access to full information about the environment and the actions of the other player. This variation abstracts away the interdependency of agent actions, meaning the agents’ actions do not directly influence each other. This allows us to focus on how humans and AI reason about each other’s goals and intentions. In the second variation, we consider a more complex and realistic situation where both players are in the same grid world environment and may collide with each other. We obtained qualitatively similar results in both variations. Due to space constraints, we have included the setup and results of the second variation in the main paper and moved the results of the first variation to the appendix. More specifically, as shown in Figure 2, the grid world is 8 by 8 in size and contains the positions of both players and four possible goals. The players can choose to move {Up, Down, Right, Left} or stay in their current grid. They can see each other’s positions and take actions simultaneously. Each player can navigate to two of the four goals. In particular, the human player can only reach one of the two goals colored blue, and the AI player can only reach one of the goals colored green. When both agents reach the same type of goal (e.g., both reaching a ”star” or both reaching a ”triangle”), they earn positive points. However, they do not earn points if they reach different types of goals or if they collide with each other (move into the same position). We set the maximum number of actions to be 20.
Figure 2. The decision-making environment in simulations and human-subject experiments. The human-AI team will receive a reward by reaching the same type of goals (star or triangle) within the time limit. They will receive no reward if they collide with each other or reach different types of goals.
4.3. Experiment 1: Evaluating Behavior Models
@@
In our first experiment, our goal is to examine the effectiveness of utilizing behavioral cloning to model human behavior. Note that his approach has been proposed and examined in various domains Carroll et al. (2019). Our main purpose to conduct this experiment is to ensure whether this method works well in our setting, as our proposed approaches in developing human belief models and collaborative AI are built on top of models of human behavior.
Experiment setup. We recruited 190 workers from Amazon Mechanical Turk. Each recruited worker was asked to play 30 navigation games within a grid world with a predefined AI model, as shown in Figure 2. To collect diverse data, we randomly paired recruited workers with one of three AI models: a random AI (selecting a random goal and taking the shortest path), a self-play AI (trained using self-play), and an optimal AI (optimized via joint optimization).
In this experiment, our goal is to develop a model of human behavior that outputs a sequence of actions given a goal. Therefore, in Experiment 1, we provided participants with additional instructions about which goal they should reach in each round, and the AI models controlled the other player to move towards the same type of goals. In our later Experiment 3, we did not provide these instructions, and humans needed to infer the goal of the AI agent.
@@
Table 1. The prediction accuracy for human behavior was evaluated assuming optimal behavior and using a data-driven model in Experiment 1. The results show that human behavior deviates significantly from optimality, and our data-driven model provides a much more accurate prediction of human behavior.
@@
Our results suggest that human behavior is significantly away from being optimal. Even when human participants are provided with suggested goals during data collection, there is only a $55.0\%$ chance that both players will reach the same type of goals across all treatments. Besides, there is a notable chance of two players colliding (about $15.8\%$) or ending in different types of goals (about $25.6\%$). We compared the performance of our learned model to a model predicated on optimal agent behavior, defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly reveal that human behavior deviates significantly from the assumed optimality. This deviation highlights the importance of incorporating a realistic model of human behavior into human-AI cooperation frameworks.
4.4. Experiment 2: Evaluating Belief Models
We next evaluate our model of human beliefs. In particular, we examine whether an AI policy that incorporates human behavior models can produce actions that are more explicable to humans, i.e., easier for humans to infer AI intentions based on AI actions. [^4] We have also directly tested the accuracy of our belief model in inferring human beliefs by comparing its predictions with direct solicitations of users’ inferences about others’ behavior. The results, included in Appendix A.3, suggest that our model is more accurate in predicting human beliefs compared to the baselines. However, they also demonstrate that human beliefs are generally quite noisy—if we randomly draw a behavior trace from another agent and ask humans to predict the goal, human prediction accuracy is close to random guessing. Therefore, we shift our focus in this experiment to whether AI agents can design their action plans to make it easier for humans to infer the goal.
Experiment setup. In this experiment, each recruited participant is presented with the behavioral traces from both agents in the environment, as shown in Figure 3. Each behavioral trace is a sequence of actions generated by an AI policy for a player. Participants are then asked to infer the intention of each agent, i.e., the goal each agent is trying to reach. If participants can better infer the goal from the behavior generated by a certain AI policy, it indicates that the AI policy is more explicable in that it generates actions that make it easier for humans to infer AI intentions.
Figure 3. The experiment interface in Experiment 2. Participants are given the behavioral traces of the agents and are asked to infer the goal each agent aims to reach.
AI policy. We now describe the set of AI policies that are used for the evaluations. As baselines, the first two AI policies are not intended to be explicable.
@@ 4.4. Experiment 2: Evaluating Belief Models
    Inexplicable-Optimal-L0: This AI policy that takes the shortest path towards the goal.
@@
    Inexplicable-Behavioral-L0: This AI policy takes actions that follow the data-driven behavioral model learned in experiment 1.
We then include two AI policies that take actions intended to be explicable. Note that to design AI policies to be explicable, we need to account for human belief models in the design of AI policies, as the AI will try to take actions that make it easier for humans to infer the goal, assuming models of human beliefs that describe how humans make inferences. The two belief models we account for are: (1) the classical level-$1$ model, where the human assumes the other agent is taking optimal level-00 actions, and (2) the behavioral level-$1$ model, where the human assumes the other agent is taking behavioral level-00 actions as described in Section 3.2. Coupled with these two belief models, and using the methodology described in Section 3.3, we can train the following two explicable AI policies.
@@
Figure 4. The results of the human-subject experiments in Experiment 2 show that when using our approach to design explicable AI, the accuracy for humans to infer AI intentions is higher than with other AI policies. Moreover, when coupled with our belief model, it results in an AI policy that makes it the easiest for humans to infer AI intentions.
@@
The experiment results are shown in Figure 4. First, we observe that when the AI is not attempting to be explicable, humans generally struggle to infer AI intentions, with the accuracy of correctly inferring the goal for both inexplicable policies being around 55% to 60%, only slightly better than random guessing. However, if we use our methodology from Section 3.3 to train the AI policy to be explicable, it significantly improves the accuracy of humans in inferring AI intentions. Moreover, when the training is coupled with our belief models, the accuracy for humans to infer AI intentions is the highest, exceeding 70%.
4.5. Experiment 3: Evaluating AI Design
We now examine our design of collaborative AI agents. Our goal is to demonstrate that the collaborative AI design that accounts for human beliefs could lead to improved human-AI collaborative performance. We utilize both simulations and human-subject experiments to examine the effectiveness of our design.
Table 2. Simulation results of collaborative performance over 10,000 testing cases in Experiment 3. The column players represent different AI agents, and the row players represent different simulated human models. The results demonstrate that while the zero-shot methods (FCP and MEP) show consistent and stable performance, when human agents possess belief models and infer AI intentions to act accordingly, designing AI that accounts for human beliefs brings significant performance improvements.
@@ 4.5. Experiment 3: Evaluating AI Design
Experiment setup. The setup of our Experiment 3 is the same as Experiment 1, shown in Figure 2. But human participants will not receive any hint about which goal their teammate or themselves are suggested to reach.
  • $\bullet$
  • $\bullet$
    Behavior-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-00 agent.
  • $\bullet$
    Belief-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-$1$ agent, i.e., it accounts for human belief models into the design of collaborative AI.
  • $\bullet$
    FCP: Fictitious Co-Play is a variant of fictitious play proposed by Strouse et al. (2021). It consists of two stages: in the first stage, a population of agents is trained independently using different random seeds and checkpoints; in the second stage, a model is trained as the best response to the agent population. In our experiment, we use the model trained with 8 partners.
  • $\bullet$
    MEP: Maximum Entropy Population-based training is proposed by  Zhao et al. (2023). The agent population is trained using population entropy bonus to increase the diversity of agents, then a single best response model is trained with respect to the agent population. In our experiment, we use the model trained with 8 partners.
We partner the collaborative AI agents with the three simulated human agents and measure the collaborative performance. The simulation results are shown in Table 2. For both the FCP and MEP baselines, despite the absence of human data in their training processes, they outperform the self-play AI when paired with the human behavior model, demonstrating their adaptability to working with teammates exhibiting unknown behavior. However, we also observe that incorporating the appropriate human behavior model directly into the training process can further enhance human-AI collaborative performance, as demonstrated by the superior performance of Behavior-AI when paired with the behavior model. Finally, when working with a behavioral level-1 agent that attempts to infer AI intentions, our proposed Belief-AI stands out, significantly outperforming other AI agents, which perform similarly to each other.
@@
These results highlight the importance of selecting appropriate human models for training collaborative AI. Furthermore, under conditions where humans attempt to infer AI intentions and act accordingly, designing AI that accounts for human beliefs can significantly improve human-AI collaborative performance.
Human-subject experiments. While the simulation results are promising, the significant performance improvement from incorporating human beliefs about AI intentions into AI design is based on the assumption that humans do indeed infer AI intentions and act accordingly. To examine whether this performance improvement carries over to collaboration with real humans, we recruited 400 workers and randomly assigned them into five treatment groups, each interacting with a different AI design. Participants were tasked with playing 30 games, preceded by three tutorial games designed to familiarize them with the gameplay. Participants could receive a \$0.05 bonus payment for reaching the same type of goal as the AI agent in each of the 30 tasks.
The experiment results are shown in Figure 5. They demonstrate that the collaborative AI trained with models of human behavior and beliefs achieved the highest collaborative performance when working with real humans. This outcome not only validates our AI design but also suggests that real human actions align with the behavior and belief models in this environment. Statistical analysis shows significant differences in the performance of AI models paired with humans, with $p<0.0001$ for comparisons between self-play AI and AI accounting for human behavior, and $p<0.0001$ for comparisons between AI accounting for human behavior versus accounting for both behavior and beliefs. The differences between MEP and FCP, as well as between FCP and behavioral-AI, are not significant ($p>0.1$). The difference between self-play AI and any other treatment is significant ($p<0.0001$). These results demonstrate that incorporating human beliefs into the design of AI agents enhances collaborative performance when working with real humans.
@@
Figure 5. The human-AI collaborative performance in our human-subject experiment in Experiment 3 shows that designing collaborative AI that accounts for human beliefs (Belief-AI treatment) significantly outperforms all other baselines when working with real humans.
5. Conclusion
Our work explores the impact of incorporating human beliefs about AI intentions into the design of collaborative AI, aiming to improve collaboration between humans and AI. By developing models that account for human beliefs regarding AI intentions and integrating these models into AI design, we have observed improved performance in human-AI collaboration. Our approach, validated through human-subject experiments, demonstrates that AI agents designed with an understanding of human behavior and beliefs over AI intention can be more effective in working with humans. This suggests a potential path forward for creating AI systems that are better aligned with human partners, making collaborative tasks more efficient and effective.
Generalization and limitations. Our approach has demonstrated effectiveness through human-subject experiments conducted in commonly used grid world environments of varying complexity. However, like most human-subject studies, our findings are limited to the chosen environments and conditions. To enhance the external validity of our approach, it is important to evaluate its performance across different domains and more realistic settings beyond grid worlds, such as dynamic, real-world applications involving physical robots or complex multi-agent systems.

### v3

@@ 3.3 Designing Explicable AI Policy
To train an explicable AI policy, we incorporate models of human beliefs into the AI training process. Rather than solely maximizing rewards, we also align the AI policy to increase the likelihood that humans can infer its intentions from its actions, using the belief model to capture the human inference process. Specifically, we use self-play to train the models. In addition to standard MDP rewards, we introduce bonus rewards when the inferred goal, based on the belief model, matches the specified goal. The bonus reward is set proportional to the log likelihood of the belief model inference, $\log(\pi(a|s,g))$. This encourages the AI policy to not only reach the goal but also to act in a way that makes its intentions more inferable, according to the belief model.
3.4 Designing Collaborative AI Agents
We now describe how we design collaborative AI agents that account for different models of human behavior and beliefs. In the experiments, we illustrate the effectiveness of incorporating models of human beliefs by examining the performance of AI agents paired with humans through simulations and real-world human-subject experiments.
@@ 3.4 Designing Collaborative AI Agents
Collaborative AI agents. We have designed several AI agents based on different assumptions of the human counterpart.
@@
    Assuming humans are optimal. We train an AI agent that learns to collaborate with itself through self-play. This is equivalent to assuming the human is acting optimally.
@@
    Assuming humans are behavioral level-00. We train an AI agent that assumes the human is a behavioral level-00 agent, using behavioral cloning to learn the human’s behavioral model.
@@
In addition to the AI agents trained using our methodologies, we also include two state-of-the-art baselines, Fictitious Co-Play (Strouse et al., 2021) and Maximum Entropy Population-based Training (Zhao et al., 2023), for comparison in our experiments. Both algorithms develop a single best response model in response to a population of agents with diverse behaviors, thus enhancing the robustness and generality of the learned best response model. The details of the setup and implementations are provided in the appendix.
4 Human-Subject Experiments
@@
To evaluate our approaches, we have conducted multiple sets of human-subject experiments through recruiting close to 1,000 participants from Amazon Mechanical Turk (MTurk). The experiments are approved by the IRB of our institution. Workers were paid \$1 base payments with the potential for bonus payments. The average hourly rate is approximately \$$\$14$. Given space constraints, we report only the key results. Additional results, including experiments in a varied domain and qualitative discussions of findings (e.g., the impact of environments), are provided in Appendix B.
4.1 Experiments Overview
@@
The main purpose of our experiments is to demonstrate that our proposed methodologies by incorporating human beliefs about AI intentions into the design of AI agents can significantly improve the performance of human-AI collaboration with real-world human participants.
(a) Experiment 1 & 3.
(b) Experiment 2.
Figure 2: The interface for our human-subject experiments. In Experiment 1, participants play with an AI agent with being told the goal, focusing on reaching a goal without colliding. In Experiment 2, participants observe agent behavior traces and infer the agent’s goal. In Experiment 3, participants decide their actions based on beliefs about AI behavior, without being told the goal.
@@
Experiment environment. Our experiments are conducted in a grid world environment with two players and multiple goals. [^3] We conducted experiments in two variations of the environment and obtained qualitatively similar results in both variations. Due to space constraints, we have included the setup and results of the second variation in the main paper and moved the results of the first variation to the appendix. Specifically, as shown in Figure 2a, the grid world is 8 by 8 in size and contains the positions of both players and four possible goals. The players can choose to move {Up, Down, Right, Left} or stay in their current grid. They can see each other’s positions and take actions simultaneously. Each player can navigate to two of the four goals: the human player can reach one of the two goals colored blue, and the AI player can reach one of the goals colored green. When both agents reach the same type of goal (e.g., both reaching a "star" or both reaching a "triangle"), they earn positive points. However, they do not earn points if they reach different types of goals or if they collide with each other (move into the same position). We set the maximum number of actions to 20.
Experiment descriptions. We conducted three sets of experiments, beginning with examining the efficacy of our models of human behavior and beliefs, and progressing to evaluate the performance improvements of adopting our design of collaborative AI. In Experiment 1, we examine the effectiveness of using behavioral cloning as a model to mimic human behavior. We recruit real-world participants to play against AI agents to collect their behavioral traces and then evaluate the prediction accuracy of using behavioral cloning to model human behavior. In Experiment 2, we examine whether we can leverage the model of human beliefs to design an explicable AI policy (as described in Section 3.3). Specifically, we recruit real-world participants, provide them with behavioral traces generated by different AI policies, and evaluate whether the AI policies that account for human beliefs produce behavioral traces that make it easier for humans to infer AI intentions. Finally, in Experiment 3, we examine whether the collaborative AI that accounts for human beliefs lead to higher performance in human-AI collaboration. Again, we recruit real-world human participants, pair them with different AI agents (as described in Section 3.4), and measure the overall team performance.
4.2 Experiment 1: Evaluating Behavior Models
@@
In our first experiment, our goal is to examine the effectiveness of utilizing behavioral cloning to model human behavior. Note that this approach has been proposed and examined in various domains (Carroll et al., 2019). This experiment is a replication of prior works in our setting. The main purpose is to examine whether this method works well in our setting, as our proposed approaches in developing human belief models and collaborative AI are built on top of models of human behavior.
Experiment setup. We recruited 190 workers from Amazon Mechanical Turk. Each participant was asked to play 30 navigation games within a grid world (as shown in Figure 2a) alongside an AI model. The goal of this experiment was to develop a model of human behavior that predicts a sequence of actions given a specified goal. To this end, we instructed participants on which goal to reach in each round and recorded their behavior as the collected dataset.
@@
Table 1: The prediction accuracy for human behavior was evaluated assuming optimal behavior and using a data-driven model in Experiment 1.
@@
We compared the performance of our learned model with a model predicated on optimal agent behavior, defined as taking the shortest path to the goal. The training, validation, and test accuracies of both models are presented in Table 1. These results clearly reveal that human behavior deviates significantly from the assumed optimality. This deviation highlights the importance of incorporating a realistic model of human behavior into human-AI cooperation frameworks. Moreover, the data-driven model trained using behavioral cloning achieves reasonable and much higher predictive performance.
4.3 Experiment 2: Evaluating Belief Models
We next evaluate our model of human beliefs. In particular, we examine whether the belief model can enable us to develop explicable AI policy (as described in Section 3.3), making it easier for humans to infer AI intention based on AI actions. [^4] We also evaluated the accuracy of our belief model by comparing its predictions with users’ directly reported inferences about others’ behavior. As shown in Appendix B.3, our model outperforms baseline predictors of human beliefs. However, the results also reveal that human beliefs are highly noisy. When shown a randomly selected behavior trace, human goal inference is often close to random guessing. Consequently, we shift the experiment’s focus to whether AI agents can plan their actions to make their goals easier for humans to infer.
Experiment setup. In this experiment, each recruited participant is presented with the behavioral traces from both agents in the environment, as shown in Figure 2b. Each behavioral trace is a sequence of actions generated by an AI policy for a player. Participants are then asked to infer the intention of each agent, i.e., the goal each agent is trying to reach. If participants can better infer the goal from the behavior generated by a certain AI policy, it indicates that the AI policy is more explicable in that it generates actions that make it easier for humans to infer AI intentions.
AI policy. We now describe the set of AI policies that are used for evaluations. As baselines, the first two AI policies are not intended to be explicable.
@@ 4.3 Experiment 2: Evaluating Belief Models
    Inexplicable-Optimal-L0: This AI takes the shortest path towards the goal.
@@
    Inexplicable-Behavioral-L0: This AI follows the behavioral model learned in Experiment 1.
We then include two AI policies that take actions intended to be explicable. To design AI policies to be explicable, we need to account for human belief models in the design of AI policies, as the AI will try to take actions that make it easier for humans to infer the goal, assuming models of human beliefs that describe how humans make inferences. The two belief models we account for are: (1) the classical level-$1$ model, where the human assumes the other agent is taking optimal level-00 actions, and (2) the behavioral level-$1$ model, where the human assumes the other agent is taking behavioral level-00 actions as described in Section 3.2. Coupled with these two belief models, and using the methodology described in Section 3.3, we train the following two explicable AI policies.
@@
Figure 3: The results of Experiment 2. Using our approach to design explicable AI leads to higher human accuracy in inferring AI intentions. Combined with our belief model, it results in an AI policy that makes it easiest to infer AI intentions.
@@
The experiment results are shown in Figure 3. First, we observe that when the AI is not attempting to be explicable, humans generally struggle to infer AI intentions, with the accuracy of correctly inferring the goal for both inexplicable policies being around 55% to 60%, only slightly better than random guessing. However, if we use our methodology from Section 3.3 to train the AI policy to be explicable, it significantly improves the accuracy of humans in inferring AI intentions. Moreover, when the training is coupled with our belief models, the accuracy for humans to infer AI intentions is much higher than all other policies, demonstrating the effectiveness of both our belief models and the methodologies in training explicable AI policy.
4.4 Experiment 3: Evaluating Collaborative AI
Finally, as the main experiment, we evaluate our collaborative AI design. The objective is to show that incorporating human beliefs into the design of collaborative AI can enhance human-AI collaborative performance through both simulations and human-subject experiments.
  • $\bullet$
  • $\bullet$
    Behavior-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-00 agent. This corresponds to the existing approaches of accounting for human behavior in AI design Carroll et al. (2019).
  • $\bullet$
    Belief-AI: the collaborative AI agent that is trained assuming the human partner is the behavioral level-$1$ agent, i.e., it accounts for human belief models into the design of AI.
  • $\bullet$
    FCP: Fictitious Co-Play is a variant of fictitious play proposed by Strouse et al. (2021). It consists of two stages: in the first stage, a population of agents is trained independently using different random seeds and checkpoints; in the second stage, a model is trained as the best response to the agent population. In our experiment, we select the model trained with 8 partners.
  • $\bullet$
    MEP: Maximum Entropy Population-based training is proposed by  Zhao et al. (2023). The agent population is trained using population entropy bonus to increase the diversity of agents, then a single best response model is trained with respect to the agent population. In our experiment, we select the model trained with 8 partners.
We emphasize again that this set of comparisons spans a wide range of state-of-the-art baselines in the literature on collaborative AI design, including standard multi-agent RL approaches (e.g., self-play), methods that account for human behavior (Carroll et al., 2019) (Behavior-AI), and zero-shot population-based approaches (FCP and MEP). Our approach is distinct in that it incorporates a model of human belief into the design of collaborative AI—to our knowledge, the first to do so.
We partner the collaborative AI agents with the three simulated human agents and measure the collaborative performance. The simulation results are shown in Table 2. For both the FCP and MEP baselines, despite the absence of human data in their training processes, they outperform the self-play AI when paired with the human behavior model, demonstrating their adaptability to working with teammates exhibiting unknown behavior. However, we also observe that incorporating the appropriate human behavior model directly into the training process can further enhance human-AI collaborative performance, as demonstrated by the superior performance of Behavior-AI when paired with the behavior model. Finally, when working with a behavioral level-1 agent that attempts to infer AI intentions, our proposed Belief-AI stands out, significantly outperforming other AI agents.
Table 2: Simulation results of collaborative performance over 10,000 testing cases in Experiment 3. The column players represent different AI agents, and the row players represent different simulated human models. The results demonstrate that while the zero-shot methods (FCP and MEP) show stable performance, when human agents infer AI intentions to act accordingly, designing AI that accounts for human beliefs brings significant performance improvements.
@@ 4.4 Experiment 3: Evaluating Collaborative AI
These results highlight the importance of incorporating appropriate human models (of behavior and beliefs) for training collaborative AI. Furthermore, under conditions where humans attempt to infer AI intentions and act accordingly, designing AI that accounts for human beliefs can significantly improve human-AI collaborative performance.
@@
Figure 4: The human-AI collaborative performance in Experiment 3. Designing collaborative AI that accounts for human beliefs (Belief-AI treatment) significantly outperforms all other baselines when paired with real humans.
Human-subject experiments. While the simulation results are promising, the significant performance improvement from incorporating human beliefs into AI design is based on the assumption that humans follow our belief model. To examine whether this performance improvement carries over to collaborations with real humans, we recruited 400 workers and randomly assigned them into five treatment groups, each interacting with a different AI design. Participants were tasked with playing 30 games, preceded by three tutorial games designed to familiarize them with the gameplay. Participants could receive a \$0.05 bonus payment for reaching the same type of goal as the AI agent in each of the 30 tasks.
We first examine whether the data-driven model learned in Experiment 1 still makes good prediction on human behavior in Experiment 3 (i.e., whether the learned behavior model performance generalizeds to new experiment). Overall, we find that the model achieves a accuracy of $73.19\%$ in predicting human behavior in Experiment 3 across all treatments, demonstrating the generalizability of learned human behavior model. More detailed discussion is provided in Appendix B.5.
We then examine the collaborative performance for pairing human participants with different AI design. The experiment results are shown in Figure 4. They demonstrate that the collaborative AI trained with models of human beliefs achieved the highest collaborative performance when working with real humans, significantly outperforming other AI designs. This outcome not only validates our AI design but also suggests that real human actions align with our belief models in this environment.
More formally, statistical analysis shows significant differences in the performance of AI models paired with humans, with $p<0.0001$ for comparisons between self-play AI and Behavior-AI, and $p<0.0001$ for comparisons between Behavior-AIversus Belief-AI. The differences between MEP and FCP, as well as between FCP and Behavior-AI, are not significant ($p>0.1$). The difference between self-play AI and any other treatments is significant ($p<0.0001$).
5 Conclusion and Discussion
Our work investigates whether accounting for human beliefs about AI intentions can improve human-AI collaboration. We make several key contributions. First, we extend level-k frameworks to develop human belief models that are empirically more accurate. Second, we introduce general methodologies for training collaborative AI agents that incorporate these models—the first to do so, to our knowledge. Finally, through extensive human-subject experiments, we show that AI agents designed with an understanding of human beliefs about AI intentions lead to improved collaborative performance with real-world users. This underscores the importance of considering human beliefs in the design of collaborative AI and demonstrates the effectiveness of both our belief model and agent design. Overall, our work points to a promising direction for building AI systems that more effectively align with their human partners.